OpenStudy (anonymous):

Need help will give medal!! We are doing systems of equations in three variables... -3a - 4b + 2c = 28 a +3b - 4c = -31 2a + 3c = 11

OpenStudy (anonymous):

Will you help? @texaschic101

OpenStudy (texaschic101):

-3a - 4b + 2c = 28 ---(multiply by 3) a + 3b - 4c = -31----(multiply by 4) -------------- -9a - 12b + 6c = 84 (result of multiplying by 3) 4a + 12b - 16c = - 124 (result of multiplying by 4) --------------add -5a - 10c = - 40 2a + 3c = 11 ----(multiply by 5) -5a - 10c = -40 ---(multiply by 2) ------------- 10a + 15c = 55 (result of multiplying by 5) -10a - 20c = -80 (result of multiplying by 2) --------------add -5c = - 25 c = 5 2a + 3c = 11 2a + 3(5) = 11 2a + 15 = 11 2a = 11 - 15 2a = -4 a = -2 a + 3b - 4c = -31 -2 + 3b - 4(5) = -31 -2 + 3b - 20 = -31 -22 + 3b = -31 3b = -31 + 22 3b = - 9 b = -3 check... -3a - 4b + 2c = 28 -3(-2) - 4(-3) + 2(5) = 28 6 + 12 + 10 = 28 28 = 28 (correct) so...a = -2, b = -3, and c = 5

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!! I have one more question if you don't mind helping me

OpenStudy (texaschic101):

I don't mind...but I can't make any promises that I will know it

OpenStudy (anonymous):

its ok you are 100 times smarter than me and i appreciate your help

OpenStudy (anonymous):

4x + 2y + 6z = 61 -12x + 3y - 5z + 8 -4x + 7y + 7z + 34

OpenStudy (texaschic101):

4x + 2y + 6z = 61---(multiply by 3) -12x + 3y - 5z = 8 ----------------- 12x + 6y + 18z = 183 (result of multiplying by 3) -12x + 3y - 5z = 8 ----------------add 9y + 13z = 175 4x + 2y + 6z = 61 -4x + 7y + 7z = 34 ----------------add 9y + 13z = 95 9y + 13z = 175 -- multiply by -1 9y + 13z = 95 ---------------- -9y - 13z = -175 9y + 13z = 95 ------------------add 0 = - 80 (INCORRECT) This has no solution

OpenStudy (texaschic101):

sorry it took so long but I was having problems with it....I do not think it has a solution. I am sorry but I have to go now....I will be back on later.

OpenStudy (anonymous):

Thank you again!!!!

OpenStudy (anonymous):

x + 2y + 12 3y - 4z = 25 x + 6y + z = 20 will you help me with this one??

OpenStudy (anonymous):

Will you please help me?? @Directrix

Directrix (directrix):

Something is missing from the first equation you posted. x + 2y + 12 equals what?

Directrix (directrix):

@~Joelie~ ^^^^

OpenStudy (anonymous):

oh my bad its suppose to be x + 2y = 12 not + 12

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Click on the attachments.